Define a new operation A*B = (A + B) • B. For example, 2 * 3 = 15. Evaluate (3*5)*7.
Answer by jessica43(140)

So order of operations tells us that we need to perform the operation in parentheses first. Se we will calculate (3*5).
A * B = (A + B) • B
3 * 5 = (3 + 5) • 5
3 * 5 = (8) • 5
3 * 5 = 40
Now we plug in 40 into the orginal equation and then follow the same steps:
(3*5)*7
(40)*7
A * B = (A + B) • B
40 * 7 = (40 + 7) • 7
40 * 7 = (47) • 7
40 * 7 = 329
So (3*5)*7 = 329